# MicroLad: 2D-to-3D Microstructure Reconstruction and Generation via Latent Diffusion and Score Distillation

### v3

@@ 5.3 Slice connectivity and spatial coherence
In the meantime, an important point of discussion is that the present 2D-to-3D microstructure reconstruction/generation framework relies on the assumption that spatial connectivity in 3D space can be represented by connectivity along three orthogonal directions (Eq. (5)). The local consistency assumption (Section 3.4) was likewise built on the optimization of latent slices along these three directions. Although LDM with a microstructure encoder (i.e., VAE) was adopted to enable the computationally efficient 2D-to-3D expansion and to generate 2D slices that remain close to one another in the latent embedding space (Table 3 and Figure 10), consistency in other directions in 3D space (e.g., diagonal) [66] is not explicitly guaranteed. While the proposed approach and its underlying assumptions are practical, given that most experimentally observed 2D micrographs are obtained along one axis or a few orthogonal planes, statistical consistency in non-orthogonal directions or planes in 3D space also needs to be evaluated and addressed in future work. Nevertheless, the proposed approach performed well for both binary and three-phase microstructures, and additional examples, including anisotropic cases presented in Appendix G, further demonstrate that the proposed framework can effectively handle a various types of microstructures.
@@ 5.4 Computational cost
Since the proposed MicroLad framework employs latent diffusion in a lower-dimensional space, the L-MPDD procedure for generating initial 3D microstructures is highly efficient. Generating a $64\times 64\times 64$ volume takes less than 10 seconds—comparable to generating a single $64\times 64$ image with a batch size of 64 (wall-clock times were measured on a single NVIDIA RTX 4090 GPU). In contrast, the original MPDD framework required over 30 minutes to produce a single 3D sample with comparable accuracy [51]. While dimensionality expansion via L-MPDD is fast, the most computationally intensive component is SDS applied to enforce target objectives. For each objective, the wall-clock time increases approximately linearly with the number of SDS steps. Appendix H presents a parametric study evaluating the effects of SDS learning rate and objective weights on optimization performance. For volume fraction and surface area targets, 500 SDS steps typically require about one minute to generate a single 3D sample. In contrast, SDS targeting effective diffusivity requires approximately two and a half minutes per sample, due to the overhead of computing gradients through the differentiable FEM solver. Across all objectives, using 3000 SDS steps consistently results in low error between generated samples and their respective targets. Thus, the number of SDS steps can be chosen by balancing computational cost with the desired accuracy of inverse generation.
@@
It is also worth noting that the main purpose of adopting a differentiable model for guided generation is to eliminate the need for pre-labeling data and training conditional models by directly integrating differentiable simulations into the guided generation framework. This formulation can be adapted depending on the application. If a sufficiently large dataset is available (or can be affordably generated) and reducing inference time is critical—perhaps even more important than minimizing training cost—a conditional formulation (e.g., conditional diffusion models with classifier-free guidance [67]) may be more suitable. In such cases, differentiable model–based guidance may no longer be necessary. The limitation of this approach, however, is that the achievable design or property range tends to be restricted by the data-conditioned generation process, whereas guided generation with differentiable models emphasizes optimization while leveraging the diffusion prior to ensure that generation remains within the unconditional data manifold [68, 53]. Therefore, the appropriate formulation should be selected based on the user’s specific objectives and design requirements.
@@
In addition, to further illustrate the computational cost and scalability of the proposed approach, we provided a 2D-to-3D reconstruction and generation case with a larger microstructure size ($256\times 256\times 256$) in Appendix I. The reconstruction of a 3D volume of this size required only 71 s, with a maximum two-point correlation function relative error of 5.23%, demonstrating both the efficiency and accuracy of the proposed framework for larger microstructure size. Meanwhile, the $D_{e}$-guided generation successfully shifted the property distribution toward the target but required 156.1 minutes, primarily because the FEM computation substantially increased the overall computational cost. This efficiency can be further improved by employing more efficient FEM solvers and differentiable neural network–based surrogate modeling approaches [69, 70], which would enable faster and more efficient gradient computation. Incorporating such surrogates represents a promising direction for future work.
@@ 5.5 Future work
The proposed MicroLad framework presents considerable potential for extension to a broader class of multiphase materials, enabling the study of microstructural variability and its influence on resulting material properties. One promising direction is to leverage the framework for uncertainty quantification by incorporating user-defined microstructural variability, thereby enhancing its utility in data-driven material design. Furthermore, MicroLad can be extended to a wider range of physics-informed generative modeling tasks, including 2D-to-3D microstructure generation under multiphysics loading conditions and complex material behavior. A particularly promising avenue for future research is the integration of advanced differentiable solvers [69] capable of modeling nonlinear constitutive laws. Such an extension would allow the framework to capture a richer spectrum of microstructural phenomena governed by diverse physical mechanisms. The incorporation of a wider variety of microstructures, including those with more phases and even continuous-valued features (e.g., crystallographic orientations), is also suggested, as the proposed approach does not intrinsically restrict the input to discrete phases. This is supported by the fact that latent diffusion models have been successfully applied to colored images as well [48]
On the generative modeling side, the current implementation of MicroLad applies SDS to enable property- and descriptor-guided generation without requiring pre-labeled microstructure–descriptor or microstructure–property pairs. This approach is particularly advantageous in scenarios where labeled data are scarce due to computational or experimental constraints. However, a natural extension would be to train conditional diffusion models that accept target objectives as direct inputs through conditioning mechanisms [67, 41]. This would eliminate the need for iterative SDS optimization and enable more efficient and scalable inverse-controlled generation.
@@ Appendix A Refinement of 2D slices on orthogonal planes
To handle unexpected noise or deviations that may occur during the 2D-to-3D dimensionality expansion, a refinement step is applied to the reconstructed 3D volume. This process does not involve training a new denoising model. Instead, it utilizes the originally trained VAE to refine the 2D slices. The key idea is that a VAE trained exclusively on clean data implicitly projects inputs back onto the learned manifold. Because the latent space captures only the distribution of clean samples, details inconsistent with the training distribution are not faithfully reconstructed [71, 72, 73]. Therefore, for the generated 3D volumes, this serves as an additional refinement step applied to the slices along the three orthogonal directions, as illustrated in Figure A.1.
@@ Appendix B Configuration parameters for the models
After training the VAE, we train a LDM that operates on the compressed latent space to enable generative modeling. Following the DDPM formulation [42], the LDM defines a forward Markov diffusion process that gradually adds Gaussian noise to a clean latent code. The LDM is trained with 1000 diffusion steps using a linear noise schedule $\beta\in[1\times 10^{-4},2\times 10^{-2}]$. During inference, new samples are generated by reversing the diffusion process starting from pure noise in latent space. These are decoded back to pixel space using the pretrained VAE decoder $\mathcal{D}(\cdot)$. The training hyperparameters for both the VAE and LDM are summarized in Tables B.2 and B.3. The encoder–decoder architecture and full latent U-Net pipeline are detailed in Table B.4,
@@ Appendix F Latent-space embedding and UMAP projection
To visualize and analyze the structure of generated microstructures in latent space, we project a collection of 2D slices from 3D volumes onto a two-dimensional manifold using UMAP [62]. The process begins by passing each microstructure slice $\mathbf{x}$ through the pretrained encoder $E_{\boldsymbol{\theta}}$, which produces spatial mean and log-variance maps:
@@ Appendix G Additional microstructure reconstruction results including anisotropic cases
To demonstrate the capability of the proposed framework in reconstructing diverse types of microstructures, several additional examples were considered, including spherical particle inclusion microstructures [51] and anisotropic battery separators [74]. As shown in Figure G.1(a), the $64\times 64$ images of spherical particle inclusions were used to train the LDM, and the MicroLad framework was then applied to perform 2D-to-3D reconstruction. The reconstructed 3D microstructures are presented in Figure G.1(b), and their corresponding 2D slices are shown in Figure G.1(c). Although some reconstructed particles appear slightly deformed (i.e., not perfectly spherical) in the 2D slices, the overall 3D spherical morphology is well reproduced, as seen in Figure G.1(b). Furthermore, the two-point correlation functions of the reconstructed samples show very good agreement with those of the original dataset (Figure G.1(d)), with a relative error of only 0.7% under $S_{2}$ guidance.
@@
In order to further demonstrate the capability of the proposed MicroLad framework for anisotropic microstructures, a conditional 2D LDM was trained to generate 2D microstructures corresponding to the xz, xy, and yz planes of a battery separator, similar to the anisotropic microstructure reconstruction version of MPDD presented in [52]. A single image was prepared for each plane: xz ($650\times 1300$), xy ($650\times 499$), and yz ($499\times 1300$). The cropped $64\times 64$ regions from these images were used to train the conditional 2D LDM to generate samples based on the specified plane orientation. The MicroLad framework was then applied to reconstruct the 3D anisotropic microstructures by using the conditional LDM to denoise the 2D latent slices corresponding to each plane. Figure G.2(a) shows the reconstructed 3D anisotropic microstructures, which were obtained using a 2D LDM trained on the corresponding 2D dataset (Figure G.2(b)). The reconstructed 3D volumes and their 2D slices exhibit strong visual similarity across different planes, indicating successful reconstruction of anisotropic features. Moreover, the two-point correlation function comparisons for each plane (Figures G.2(d)–(f)) demonstrate that the spatial correlations were well reproduced across the planes, with a maximum relative error of 2.6% (on the yz plane). These results highlight the effectiveness of the proposed framework for 2D-to-3D reconstruction of anisotropic microstructures.
@@
It is worth noting that, until now, the microstructures considered in this study are not perfectly isotropic or orthogonally anisotropic (i.e., they exhibit certain anisotropic features in non-orthogonal directions). Nevertheless, the visual similarities and spatial correlations were well preserved. To further demonstrate that the proposed MicroLad framework can reconstruct microstructures with pronounced non-orthogonal anisotropy, a battery separator microstructure rotated by $45^{\circ}$ on the xz-plane was used for 2D-to-3D reconstruction, as shown in Figure G.3. As illustrated in Figure G.3(a), the 3D microstructures were reconstructed based on a dataset exhibiting diagonal anisotropy (Figure G.3(b)), and the 2D slices of the reconstructed 3D volumes (Figure G.3(c)) confirm that the non-orthogonal anisotropy was successfully reproduced. The two-point correlation function comparisons (Figures G.3(d)–(e)) show a maximum relative error of only 1.4% on the yz-plane, further highlighting that the spatial correlations were well preserved. These results demonstrate the proposed framework’s capability to handle non-orthogonal anisotropy effectively. In addition, Bostanabad [66] discussed that non-orthogonal anisotropy can be effectively captured in orthogonality-based microstructure reconstruction when the voxel field is updated simultaneously rather than through individual voxel updates. In this context, the performance of the proposed framework can be attributed to the MPDD mechanism, which performs multi-plane denoising processes across a large number of diffusion steps. This process effectively approximates a continuous stochastic differential equation, as discussed in [51], allowing the entire voxel field—rather than individual voxels—to be updated simultaneously from the prior to the data distribution across multiple planes.
@@
Figure G.2: 2D-to-3D microstructure reconstruction results for anisotropic microstructures (lithium-ion battery separator [74]): (a) reconstructed 3D microstructure, (b) original 2D microstructure images, (c) random cross-sectional slices of (a), and (d)–(f) two-point correlation functions of 2D slices from the reconstructed 3D samples in (a) compared to those from the original 2D dataset in (b) for the xz, xy, and yz planes, respectively.
@@ Appendix I Experiments with larger microstructure size
Next, to validate the inverse-controlled 2D-to-3D generation performance for larger microstructure sizes, the inverse generation results targeting $V_{f}$, $SA$, and $D_{e}$ are illustrated in Figures I.2, I.3, and I.4, respectively. As shown in these figures, the proposed MicroLad framework successfully generates microstructures that reach the maximum values of the dataset for each objective (i.e., the tail of the data distribution), consistent with the $64\times 64\times 64$ cases presented in Section 4.2. These results highlight the capability of the proposed framework for data expansion and inverse-controlled generation at larger microstructural scales. The inverse-controlled generation used 9000 SDS steps, requiring 21.6 minutes for volume-fraction control and 23.6 minutes for surface-area control to generate a single 3D volume. For $D_{e}$-guided generation, the process took 156.1 minutes because the FEM computations for $256\times 256$ microstructure images substantially increased the overall computational cost compared to the $64\times 64$ case, as expected. Nonetheless, this computational efficiency can be further improved by employing more efficient FEM solvers and surrogate modeling approaches [69, 70], which enable differentiable and near-instantaneous evaluation of $D_{e}$. Such improvements are suggested as promising directions for future work.

### v4

@@ 5.3 Slice connectivity and spatial coherence
In the meantime, an important point of discussion is that the present 2D-to-3D microstructure reconstruction/generation framework relies on the assumption that spatial connectivity in 3D space can be represented by connectivity along three orthogonal directions (Eq. (5)). The local consistency assumption (Section 3.4) was likewise built on the optimization of latent slices along these three directions. Although LDM with a microstructure encoder (i.e., VAE) was adopted to enable the computationally efficient 2D-to-3D expansion and to generate 2D slices that remain close to one another in the latent embedding space (Table 3 and Figure 10), consistency in other directions in 3D space (e.g., diagonal) [Bostanabad2020Reconstruction] is not explicitly guaranteed. While the proposed approach and its underlying assumptions are practical, given that most experimentally observed 2D micrographs are obtained along one axis or a few orthogonal planes, statistical consistency in non-orthogonal directions or planes in 3D space also needs to be evaluated and addressed in future work. Nevertheless, the proposed approach performed well for both binary and three-phase microstructures, and additional examples, including anisotropic cases presented in Appendix G, further demonstrate that the proposed framework can effectively handle a various types of microstructures.
@@ 5.4 Computational cost
Since the proposed MicroLad framework employs latent diffusion in a lower-dimensional space, the L-MPDD procedure for generating initial 3D microstructures is highly efficient. Generating a $64\times 64\times 64$ volume takes less than 10 seconds—comparable to generating a single $64\times 64$ image with a batch size of 64 (wall-clock times were measured on a single NVIDIA RTX 4090 GPU). In contrast, the original MPDD framework required over 30 minutes to produce a single 3D sample with comparable accuracy [lee2024multi]. While dimensionality expansion via L-MPDD is fast, the most computationally intensive component is SDS applied to enforce target objectives. For each objective, the wall-clock time increases approximately linearly with the number of SDS steps. Appendix H presents a parametric study evaluating the effects of SDS learning rate and objective weights on optimization performance. For volume fraction and surface area targets, 500 SDS steps typically require about one minute to generate a single 3D sample. In contrast, SDS targeting effective diffusivity requires approximately two and a half minutes per sample, due to the overhead of computing gradients through the differentiable FEM solver. Across all objectives, using 3000 SDS steps consistently results in low error between generated samples and their respective targets. Thus, the number of SDS steps can be chosen by balancing computational cost with the desired accuracy of inverse generation.
@@
It is also worth noting that the main purpose of adopting a differentiable model for guided generation is to eliminate the need for pre-labeling data and training conditional models by directly integrating differentiable simulations into the guided generation framework. This formulation can be adapted depending on the application. If a sufficiently large dataset is available (or can be affordably generated) and reducing inference time is critical—perhaps even more important than minimizing training cost—a conditional formulation (e.g., conditional diffusion models with classifier-free guidance [ho2022classifier]) may be more suitable. In such cases, differentiable model–based guidance may no longer be necessary. The limitation of this approach, however, is that the achievable design or property range tends to be restricted by the data-conditioned generation process, whereas guided generation with differentiable models emphasizes optimization while leveraging the diffusion prior to ensure that generation remains within the unconditional data manifold [maze2023diffusion, poole2022dreamfusion]. Therefore, the appropriate formulation should be selected based on the user’s specific objectives and design requirements.
@@
In addition, to further illustrate the computational cost and scalability of the proposed approach, we provided a 2D-to-3D reconstruction and generation case with a larger microstructure size ($256\times 256\times 256$) in Appendix I. The reconstruction of a 3D volume of this size required only 71 s, with a maximum two-point correlation function relative error of 5.23%, demonstrating both the efficiency and accuracy of the proposed framework for larger microstructure size. Meanwhile, the $D_{e}$-guided generation successfully shifted the property distribution toward the target but required 156.1 minutes, primarily because the FEM computation substantially increased the overall computational cost. This efficiency can be further improved by employing more efficient FEM solvers and differentiable neural network–based surrogate modeling approaches [xue2023jax, kudela2022recent], which would enable faster and more efficient gradient computation. Incorporating such surrogates represents a promising direction for future work.
@@ 5.5 Future work
The proposed MicroLad framework presents considerable potential for extension to a broader class of multiphase materials, enabling the study of microstructural variability and its influence on resulting material properties. One promising direction is to leverage the framework for uncertainty quantification by incorporating user-defined microstructural variability, thereby enhancing its utility in data-driven material design. Furthermore, MicroLad can be extended to a wider range of physics-informed generative modeling tasks, including 2D-to-3D microstructure generation under multiphysics loading conditions and complex material behavior. A particularly promising avenue for future research is the integration of advanced differentiable solvers [xue2023jax] capable of modeling nonlinear constitutive laws. Such an extension would allow the framework to capture a richer spectrum of microstructural phenomena governed by diverse physical mechanisms. The incorporation of a wider variety of microstructures, including those with more phases and even continuous-valued features (e.g., crystallographic orientations), is also suggested, as the proposed approach does not intrinsically restrict the input to discrete phases. This is supported by the fact that latent diffusion models have been successfully applied to colored images as well [rombach2022high]
On the generative modeling side, the current implementation of MicroLad applies SDS to enable property- and descriptor-guided generation without requiring pre-labeled microstructure–descriptor or microstructure–property pairs. This approach is particularly advantageous in scenarios where labeled data are scarce due to computational or experimental constraints. However, a natural extension would be to train conditional diffusion models that accept target objectives as direct inputs through conditioning mechanisms [ho2022classifier, yang2023diffusion]. This would eliminate the need for iterative SDS optimization and enable more efficient and scalable inverse-controlled generation.
@@ Appendix A Refinement of 2D slices on orthogonal planes
To handle unexpected noise or deviations that may occur during the 2D-to-3D dimensionality expansion, a refinement step is applied to the reconstructed 3D volume. This process does not involve training a new denoising model. Instead, it utilizes the originally trained VAE to refine the 2D slices. The key idea is that a VAE trained exclusively on clean data implicitly projects inputs back onto the learned manifold. Because the latent space captures only the distribution of clean samples, details inconsistent with the training distribution are not faithfully reconstructed [hassanaly2024evaluation, meng2017magnet, chen2018unsupervised]. Therefore, for the generated 3D volumes, this serves as an additional refinement step applied to the slices along the three orthogonal directions, as illustrated in Figure A.1.
@@ Appendix B Configuration parameters for the models
After training the VAE, we train a LDM that operates on the compressed latent space to enable generative modeling. Following the DDPM formulation [ho2020denoising], the LDM defines a forward Markov diffusion process that gradually adds Gaussian noise to a clean latent code. The LDM is trained with 1000 diffusion steps using a linear noise schedule $\beta\in[1\times 10^{-4},2\times 10^{-2}]$. During inference, new samples are generated by reversing the diffusion process starting from pure noise in latent space. These are decoded back to pixel space using the pretrained VAE decoder $\mathcal{D}(\cdot)$. The training hyperparameters for both the VAE and LDM are summarized in Tables B.2 and B.3. The encoder–decoder architecture and full latent U-Net pipeline are detailed in Table B.4,
@@ Appendix F Latent-space embedding and UMAP projection
To visualize and analyze the structure of generated microstructures in latent space, we project a collection of 2D slices from 3D volumes onto a two-dimensional manifold using UMAP [mcinnes2018umap]. The process begins by passing each microstructure slice $\mathbf{x}$ through the pretrained encoder $E_{\boldsymbol{\theta}}$, which produces spatial mean and log-variance maps:
@@ Appendix G Additional microstructure reconstruction results including anisotropic cases
To demonstrate the capability of the proposed framework in reconstructing diverse types of microstructures, several additional examples were considered, including spherical particle inclusion microstructures [lee2024multi] and anisotropic battery separators [lagadec2018microstructure]. As shown in Figure G.1(a), the $64\times 64$ images of spherical particle inclusions were used to train the LDM, and the MicroLad framework was then applied to perform 2D-to-3D reconstruction. The reconstructed 3D microstructures are presented in Figure G.1(b), and their corresponding 2D slices are shown in Figure G.1(c). Although some reconstructed particles appear slightly deformed (i.e., not perfectly spherical) in the 2D slices, the overall 3D spherical morphology is well reproduced, as seen in Figure G.1(b). Furthermore, the two-point correlation functions of the reconstructed samples show very good agreement with those of the original dataset (Figure G.1(d)), with a relative error of only 0.7% under $S_{2}$ guidance.
@@
In order to further demonstrate the capability of the proposed MicroLad framework for anisotropic microstructures, a conditional 2D LDM was trained to generate 2D microstructures corresponding to the xz, xy, and yz planes of a battery separator, similar to the anisotropic microstructure reconstruction version of MPDD presented in [lee2024denoising]. A single image was prepared for each plane: xz ($650\times 1300$), xy ($650\times 499$), and yz ($499\times 1300$). The cropped $64\times 64$ regions from these images were used to train the conditional 2D LDM to generate samples based on the specified plane orientation. The MicroLad framework was then applied to reconstruct the 3D anisotropic microstructures by using the conditional LDM to denoise the 2D latent slices corresponding to each plane. Figure G.2(a) shows the reconstructed 3D anisotropic microstructures, which were obtained using a 2D LDM trained on the corresponding 2D dataset (Figure G.2(b)). The reconstructed 3D volumes and their 2D slices exhibit strong visual similarity across different planes, indicating successful reconstruction of anisotropic features. Moreover, the two-point correlation function comparisons for each plane (Figures G.2(d)–(f)) demonstrate that the spatial correlations were well reproduced across the planes, with a maximum relative error of 2.6% (on the yz plane). These results highlight the effectiveness of the proposed framework for 2D-to-3D reconstruction of anisotropic microstructures.
@@
It is worth noting that, until now, the microstructures considered in this study are not perfectly isotropic or orthogonally anisotropic (i.e., they exhibit certain anisotropic features in non-orthogonal directions). Nevertheless, the visual similarities and spatial correlations were well preserved. To further demonstrate that the proposed MicroLad framework can reconstruct microstructures with pronounced non-orthogonal anisotropy, a battery separator microstructure rotated by $45^{\circ}$ on the xz-plane was used for 2D-to-3D reconstruction, as shown in Figure G.3. As illustrated in Figure G.3(a), the 3D microstructures were reconstructed based on a dataset exhibiting diagonal anisotropy (Figure G.3(b)), and the 2D slices of the reconstructed 3D volumes (Figure G.3(c)) confirm that the non-orthogonal anisotropy was successfully reproduced. The two-point correlation function comparisons (Figures G.3(d)–(e)) show a maximum relative error of only 1.4% on the yz-plane, further highlighting that the spatial correlations were well preserved. These results demonstrate the proposed framework’s capability to handle non-orthogonal anisotropy effectively. In addition, Bostanabad [Bostanabad2020Reconstruction] discussed that non-orthogonal anisotropy can be effectively captured in orthogonality-based microstructure reconstruction when the voxel field is updated simultaneously rather than through individual voxel updates. In this context, the performance of the proposed framework can be attributed to the MPDD mechanism, which performs multi-plane denoising processes across a large number of diffusion steps. This process effectively approximates a continuous stochastic differential equation, as discussed in [lee2024multi], allowing the entire voxel field—rather than individual voxels—to be updated simultaneously from the prior to the data distribution across multiple planes.
@@
Figure G.2: 2D-to-3D microstructure reconstruction results for anisotropic microstructures (lithium-ion battery separator [lagadec2018microstructure]): (a) reconstructed 3D microstructure, (b) original 2D microstructure images, (c) random cross-sectional slices of (a), and (d)–(f) two-point correlation functions of 2D slices from the reconstructed 3D samples in (a) compared to those from the original 2D dataset in (b) for the xz, xy, and yz planes, respectively.
@@ Appendix I Experiments with larger microstructure size
Next, to validate the inverse-controlled 2D-to-3D generation performance for larger microstructure sizes, the inverse generation results targeting $V_{f}$, $SA$, and $D_{e}$ are illustrated in Figures I.2, I.3, and I.4, respectively. As shown in these figures, the proposed MicroLad framework successfully generates microstructures that reach the maximum values of the dataset for each objective (i.e., the tail of the data distribution), consistent with the $64\times 64\times 64$ cases presented in Section 4.2. These results highlight the capability of the proposed framework for data expansion and inverse-controlled generation at larger microstructural scales. The inverse-controlled generation used 9000 SDS steps, requiring 21.6 minutes for volume-fraction control and 23.6 minutes for surface-area control to generate a single 3D volume. For $D_{e}$-guided generation, the process took 156.1 minutes because the FEM computations for $256\times 256$ microstructure images substantially increased the overall computational cost compared to the $64\times 64$ case, as expected. Nonetheless, this computational efficiency can be further improved by employing more efficient FEM solvers and surrogate modeling approaches [xue2023jax, kudela2022recent], which enable differentiable and near-instantaneous evaluation of $D_{e}$. Such improvements are suggested as promising directions for future work.